OpenStudy (anonymous):

If f(x)=a^(X), which of the following statements is true? a) f(x+y)=f(x)+f(y) b) f(xy)=f(x)f(y) c) (f(x))y=f(xy) d) (f(x))^(y)=f(x^y)

OpenStudy (anonymous):

none. it will be f(x+y)=f(x)f(y)

OpenStudy (anonymous):

I see you stopped asking questions on coordinate geometry!

OpenStudy (anonymous):

lol.^ and that is not a possible answer.

OpenStudy (anonymous):

Are you sure it can't be f(x+y)=f(x)+f(y) ?

OpenStudy (anonymous):

yep..it wont be tat.

OpenStudy (anonymous):

observe tat a^(x+y)=(a^x)(a^y)

OpenStudy (anonymous):

it's f(x+y) = f(x)f(y) example: 2^5 = 2^2 * 2^3 = 32

OpenStudy (anonymous):

that's essentially what saubhik said, but with an example

OpenStudy (anonymous):

so chizzle any problems?

OpenStudy (anonymous):

Yes. None of these answers fit.. Wouldn't (f(x))^(y)=f(x^(y)) work?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

however f(x)^y=f(xy) will work. I think u got some problems in exponents. Please clear them out instead of trying for the options :) Ask me if u hav any prob:)

OpenStudy (anonymous):

What do you mean clear them out instead of trying for the options?

OpenStudy (anonymous):

I mean that u can urself answer ur doubts if u had ur basics clear:) If u are uncomfortable with exponents ask me freely!

OpenStudy (anonymous):

I'm fine with exponents. It's when it comes to all variables, functions, logarithms, and having to plug things in that I get confused.

OpenStudy (anonymous):

If u know (2^3)*(2^1)=2^4 then u automatically know a^x*a^y=a^(x+y) Hint: If u have problems with variables or find difficult to deal with them assign a value to each variable given in the question and the also in the options.

OpenStudy (anonymous):

like above u could have assigned x=1 a=2 y=2. etc.. But be careful while dealing with numbers since some numbers cant be used like taking y=1 will prove d correct while its wrong actually. So u must deal with more than 1 numbers.